# Architect of the Bits World: Masked Autoregressive Modeling for Circuit Generation Guided by Truth Table

### v1

@@ 3.4 Differentiable Architecture Search
Figure 4: Scaling behavior of CircuitAR with different model parameters on circuit generation benchmark.
Figure 5: Training with more tokens improves BitsD for CircuitAR with different model parameters.
@@
Initialization. After executing Algorithm 2, the adjacency matrix of a valid DAG $\bar{\mathcal{A}}\in\mathbb{R}^{N\times N}$ and its corresponding probability matrix $\hat{\mathcal{A}}=\bar{\mathcal{A}}\cdot\tilde{\mathcal{A}}$, where $\hat{\mathcal{A}}\in\mathbb{R}^{N\times N}$, are obtained. Using $\bar{\mathcal{A}}$, we derive the hierarchical structure $H=\{h_{1},h_{2},\dots,h_{l}\}$, where $h_{l}$ represents the node list of the $l$-th layer. The set $H$ encapsulates the layer count $l$ and the width information of each layer, which is used to initialize CircuitNN illustrated in Figure 1. For connection probabilities, since each node can only connect to nodes from preceding layers, we normalize the connection probabilities such that their summation equals 1. This yields the weights $w\in\mathbb{R}^{N_{p}}$ for possible connections, where $N_{p}$ denotes the number of nodes in the previous layer. To ensure compatibility with the Softmax function applied in CircuitNN, we initialize the logits $\hat{w}\in\mathbb{R}^{N_{p}}$ such that the Softmax output matches the normalized connection probabilities. The logits are initialized as follows:
$\hat{w}=\log(w+\epsilon)-\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}\log(w_{i}+\epsilon),$ (11)
@@ 4 Experiments
4.1 Experiment Settings
Training Datasets. We generate training datasets with around 400k circuits (average 200 gates per circuit) from the open-source datasets (Bryan, 1985; Albrecht, 2005; Amarú et al., 2015). The training dataset construction details will be illustrated in Appendix B.
Data Augmentation. We provide more details about data augmentation in Appendix C and investigate the impact of idle of NAND gates in Appendix D.
Training Process. We provide more details about the training processes of CircuitVQ and CircuitAR in Section A.1 and Section A.2.
Baseline Selection. For baseline selection, we choose CircuitNN (Google Deepmind, 2023) and T-Net (Wang et al., 2024) due to their state-of-the-art (SOTA) performance in circuit generation guided by truth tables. Additionally, several other studies (Tsaras et al., 2024; Li et al., 2024b; Zhou et al., 2024) have explored circuit generation using different paradigms. We discuss these approaches in Section 5, as they diverge from the DAS paradigm employed in this work.
Evaluation. To validate the effectiveness of our CircuitAR models, we conduct evaluations using circuits from the IWLS competition (Mishchenko et al., 2022), which include five distinct function categories: random, basic functions, Espresso (Rudell, 1985), arithmetic functions, and LogicNets (Umuroglu et al., 2020). Random circuits consist of random and decomposable Boolean functions, basic functions include majority functions and binary sorters, and arithmetic functions involve arithmetic circuits with permuted inputs and dropped outputs. Furthermore, we evaluate the BitsD for CircuitAR models with different sizes to assess their conditional circuit generation capability. This evaluation is performed on our circuit generation benchmark with 1000 circuits separate from the training dataset.
4.2 Scalability and Emergent Capability
[…]
Data Scaling. Figure 6 illustrates consistent performance gains with increased training tokens across all sizes. For the 2B model, BitsD improves by 8.13% (0.4748$\rightarrow$0.4362) when scaling from 1B to 4B tokens. Moreover, larger models exhibit superior data efficiency. Specifically, the 2B model achieves better performance (0.4362 BitsD) with 4B tokens than the 1B model (0.4644 BitsD), emphasizing the interplay between model capacity and training scale.
@@ 4.3 SOTA Circuit Generation
Critically, the 100% accuracy across all benchmarks confirms that our method maintains functional correctness while achieving these efficiency gains. These results validate our hypothesis that integrating learned structural priors with CircuitAR enables more sample-efficient circuit generation compared to basic CircuitNN (Google Deepmind, 2023) and template-driven (Wang et al., 2024) DAS approaches.
@@ 4.4 Ablation Study
Table 3: Ablation study on the initialization process with edge probability generated by CircuitAR-2B.
@@
We conducted an ablation study to evaluate the effectiveness of the probability matrix $\hat{\mathcal{A}}$ generated by CircuitAR-2B. As summarized in Table 3, the experiment results reveal that both variants achieve perfect accuracy (100%) across all benchmarks, suggesting that the initialization process does not impair the ability to generate functionally equivalent circuits. The primary distinction lies in the efficiency of the search process, quantified by the number of search steps. Table 3 underscores the significance of the initialization process, demonstrating that our CircuitAR models can produce high-quality preliminary circuit structures, which can guide the subsequent DAS process effectively.
5 Related Works
5.1 Autoregressive Modeling
The autoregressive modeling paradigm (OpenAI, 2023; Tian et al., 2024) has been widely adopted for generation tasks in language and vision domains. Built on the transformer architecture, autoregressive models are commonly implemented using causal attention mechanisms in language domains (OpenAI, 2023), which process data sequentially. However, information does not inherently require sequential processing in vision and graph generation tasks. To address this, researchers have employed bidirectional attention mechanisms for autoregressive modeling (Li et al., 2024a; Tian et al., 2024; Chang et al., 2022; Li et al., 2023b). This approach predicts the next token based on previously predicted tokens while allowing unrestricted communication between tokens, thereby relaxing the sequential constraints of traditional autoregressive methods. In this paper, we adopt masked autoregressive modeling for circuit generation, leveraging its ability to provide a global perspective and enhance the modeling of complex dependencies.
5.2 Circuit Generation
In addition to the DAS-based approaches, researchers have also explored next-gate prediction techniques inspired by LLMs for circuit generation. Circuit Transformer (Li et al., 2024b) predicts the next logic gate using a depth-first traversal and equivalence-preserving decoding. SeaDAG (Zhou et al., 2024) employs a semi-autoregressive diffusion model for DAG generation, maintaining graph structure for precise control. ShortCircuit (Tsaras et al., 2024) uses a transformer to generate Boolean expressions from truth tables via next-token prediction. However, these methods are limited by their global view, restricting circuit size and failing to reduce the search space. In contrast, our approach uses global-view masked autoregressive decoding to generate circuits while ensuring logical equivalence and significantly reducing the search space during the DAS process.
@@ 6 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Implementation Details
A.1 CircuitVQ
The training process of CircuitVQ employs a linear learning rate schedule with the AdamW optimizer set at a learning rate of $2\times 10^{-4}$, a weight decay of 0.1, and a batch size of 2048. The model is fine-tuned for 20 epochs on 8$\times$×A100 GPUs with 80G memory each. Moreover, we use the Graphormer (Ying et al., 2021) as our CircuitVQ architecture as mentioned before. Specifically, CircuitVQ comprises 6 encoder layers and 1 decoder layer. The hidden dimension and FFN intermediate dimension are 1152 and 3072, respectively. Additionally, the multi-head attention mechanism employs 32 attention heads. For the vector quantizer component, the codebook dimensionality is set to 4 to improve the codebook utilization, and the codebook size is configured to 8192.
A.2 CircuitAR
The training process of CircuitAR employs a linear learning rate schedule with the AdamW optimizer set at a learning rate of $2\times 10^{-4}$, a weight decay of 0.1, and a batch size of 4096. The model is fine-tuned for 20 epochs on 16$\times$×A100 GPUs with 80G memory each. Moreover, we use the Transformer variant of Llama (Dubey et al., 2024) as our CircuitAR architecture as mentioned before. To form different model sizes, we vary the hidden dimension, FFN intermediate dimension, number of heads and number of layers. We present the details of our CirucitAR architecture configurations in Table 4. For the rest of the hyperparameters, we keep them the same as the standard Llama model.
@@ A.2 CircuitAR
Appendix B Training Datasets
@@
To mitigate data imbalance, a constraint is imposed, limiting each truth table to at most $M$ distinct graph topologies. For truth tables with fewer than $M$ representations, logic synthesis techniques (specifically rewriting algorithms) are applied to generate functionally equivalent subcircuits with distinct topologies. This approach ensures topological diversity while maintaining functional equivalence. Finally, the training datasets with around 400000 circuits (average 200 gates per circuit) are generated using circuits from the ISCAS’85 (Bryan, 1985), IWLS’05 (Albrecht, 2005), and EPFL (Amarú et al., 2015) benchmarks. The maximum input size is set to 16, and $M$ is set to 5 during the generation process.
Appendix C Data Augmentation
@@
Appendix D Idle NAND Gates
As illustrated in Figure 7, all model variants exhibit a gradual decline in BitsD as the proportion of isolated gates increases from 0% to 45%. Larger models (2B parameters), demonstrate significantly greater robustness, maintaining BitsD values within a narrow range (0.4362 to 0.4413) across varying isolation ratios. In contrast, smaller models (300M parameters) show a more pronounced degradation, with BitsD increasing from 0.5317 to 0.5484 under the same conditions. This disparity highlights the enhanced ability of larger models to efficiently utilize NAND gates for implementing the same truth table. The consistently low BitsD observed in the 2B model (0.4412 at 45% idle ratios) underscores its practical utility in predefining search spaces for DAS, offering a notable advantage over smaller models.

### v2

@@ 3.2 CircuitVQ
3.3 CircuitAR
After completing the CircuitVQ training, we train CircuitAR using a graph autoregressive modeling paradigm as shown in Figure 3, where CircuitVQ functions as the tokenizer. Let $\bm{Y}=[y_{i}]_{i=1}^{N}$ represent the discrete latent tokens of the input circuit $\mathcal{G}$, tokenized by CircuitVQ. During the masked autoregressive training process, we sample a subset of nodes $\mathcal{V}_{s}\subset\mathcal{V}$ and replace them with a special mask token $\mathit{m}$. For the masked $\bm{Y}$, the latent token $\bar{y}_{i}$ is defined as:
Following [7] and [20], we employ a cosine mask scheduling function $\gamma(r)=\cos(0.5\pi r)$ in the sampling process. This involves uniformly sampling a ratio $r$ from the interval $[0,1]$ and then selecting $\lceil\gamma(r)\cdot N\rceil$ tokens from $\bm{Y}$ to mask uniformly. Let $\bar{\bm{Y}}=[\bar{y}_{i}]_{i=1}^{N}$ denote the output after applying the masking operation to $\bm{Y}$. The masked sequence $\bar{\bm{Y}}$ is then fed into a multi-layer transformer with bidirectional attention to predict the probabilities $p(y_{i}|\bar{\bm{Y}},\mathcal{T})$ for each $v_{i}\in\mathcal{V}_{s}$ under the condition of the truth table. The transformer is designed based on Llama models, each CircuitAR layer consists of a self-attention block, a cross-attention block, and an FFN block. Specifically, the info of the truth table is conditioned by serving $\mathcal{T}$ as the input key and value of the cross-attention block. The training loss for CircuitAR is defined as:
$\mathcal{L}_{\text{ar}}=-\sum\limits_{\mathcal{D}}\sum\limits_{v_{i}\in\mathcal{V}_{s}}\log p(y_{i}|\bar{\bm{Y}},\mathcal{T}),$ (10)
where $\mathcal{D}$ represents the set of training circuits.
Autoregressive decoding. We introduce a parallel decoding method, where tokens are generated in parallel. This approach is feasible due to the bidirectional self-attention mechanism of CircuitAR. At inference time, we begin with a blank canvas $\bar{\bm{Y}}=[\mathit{m}]^{N}$ and the decoding process of CircuitAR follows Algorithm 1. Specifically, the decoding algorithm generates a circuit in $T$ steps. At each iteration, the model predicts all tokens simultaneously but retains only the most confident predictions following the cosine schedule [7, 20]. The remaining tokens are masked and re-predicted in the next iteration. The mask ratio decreases progressively until all tokens are generated within $T$ iterations.
3.4 Differentiable Architecture Search
DAG Search. The reconstructed adjacency matrix $\tilde{\mathcal{A}}$ is a probability matrix that denotes the probabilities of connections between gates. However, $\tilde{\mathcal{A}}$ may contain cycles and identifying the optimal DAG with the highest edge probabilities is an NP-hard problem. Consequently, we employ a greedy algorithm to obtain a suboptimal DAG. As illustrated in Algorithm 2, the algorithm initializes $\bar{\mathcal{A}}\in\mathbb{R}^{N\times N}$ with edge probabilities and enforces basic structural rules: PIs have no indegree, POs have no outdegree, and self-loops are prohibited in circuit designs. Following this initialization, a depth-first search (DFS) is conducted to detect cycles in $\bar{\mathcal{A}}$. If no cycles are found, $\bar{\mathcal{A}}$ is a valid DAG, and the algorithm terminates. If a cycle is detected, the edge with the lowest probability within the cycle is identified and removed by setting the corresponding edge in $\bar{\mathcal{A}}$ to 0. This process repeats iteratively until no cycles remain. This greedy approach ensures the derivation of a valid DAG $\bar{\mathcal{A}}$ that approximates the optimal structure while preserving the acyclic property necessary for circuit design. The resulting DAG serves as a foundation for further refinement in the DAS process, ultimately generating a precise circuit that is logically equivalent to $\mathcal{T}$.
@@ 3.4 Differentiable Architecture Search
Initialization. After executing Algorithm 2, the adjacency matrix of a valid DAG $\bar{\mathcal{A}}\in\mathbb{R}^{N\times N}$ and its corresponding probability matrix $\hat{\mathcal{A}}=\bar{\mathcal{A}}\cdot\tilde{\mathcal{A}}$, where $\hat{\mathcal{A}}\in\mathbb{R}^{N\times N}$, are obtained. Using $\bar{\mathcal{A}}$, we derive the hierarchical structure $H=\{h_{1},h_{2},\dots,h_{l}\}$, where $h_{l}$ represents the node list of the $l$-th layer. The set $H$ encapsulates the layer count $l$ and the width information of each layer, which is used to initialize CircuitNN illustrated in Figure 1. For connection probabilities, since each node can only connect to nodes from preceding layers, we normalize the connection probabilities such that their summation equals 1. This yields the weights $\bm{w}\in\mathbb{R}^{N_{p}}$ for possible connections, where $N_{p}$ denotes the number of nodes in the previous layer. To ensure compatibility with the Softmax function applied in CircuitNN, we initialize the logits $\hat{\bm{w}}\in\mathbb{R}^{N_{p}}$ such that the Softmax output matches the normalized connection probabilities. The logits are initialized as follows:
$\hat{\bm{w}}=\log(\bm{w}+\epsilon)-\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}\log(\bm{w}_{i}+\epsilon),$ (11)
@@ 4 Experiments
4.1 Experiment Settings
Data Augmentation. We provide more details about data augmentation in Appendix D and investigate the impact of the idle of NAND gates in Appendix E.
Training Details. We generate a training dataset with around 400k circuits (average 200 gates per circuit) from the open-source datasets [6, 2, 3]. The training dataset construction details will be illustrated in Appendix C. We also provide more details about the training processes of CircuitVQ and CircuitAR in Section B.1 and Section B.2.
Baseline Selection. For baseline selection, we choose CircuitNN [14] and T-Net [31] due to their state-of-the-art (SOTA) performance in circuit generation guided by truth tables. Additionally, several other studies [28, 21, 36] have explored circuit generation using different paradigms. We discuss these approaches in Appendix A, as they diverge from the DAS paradigm employed in this work.
Evaluation Details. To validate the effectiveness of our CircuitAR models, we conduct evaluations using circuits from the IWLS competition [23], which include five distinct function categories: random, basic functions, Espresso [25], arithmetic functions, and LogicNets [29]. Random circuits consist of random and decomposable Boolean functions, basic functions include majority functions and binary sorters, and arithmetic functions involve arithmetic circuits with permuted inputs and dropped outputs. Furthermore, we evaluate the BitsD for CircuitAR models with different sizes to assess their conditional circuit generation capability. This evaluation is performed on our circuit generation benchmark with 1000 circuits separate from the training dataset.
Figure 4: Scaling behavior of CircuitAR with different model parameters.
Figure 5: The performance of different model sizes under a fixed compute budget.
Figure 6: Training with more tokens improves BitsD for CircuitAR.
4.2 Scalability and Emergent Capability
[…]
Data Scaling. Figure 7 illustrates consistent performance gains with increased training tokens across all sizes. For the 2B model, BitsD improves by 8.13%, 0.4748$\rightarrow$0.4362, when scaling from 1B to 4B tokens. Moreover, larger models exhibit superior data efficiency. Specifically, the 2B model achieves better performance with 4B tokens than the 1B model, emphasizing the interplay between model capacity and training scale.
@@ 4.3 SOTA Circuit Generation
Critically, the 100% accuracy across all benchmarks confirms that our method maintains functional correctness while achieving these efficiency gains. This is guaranteed by the DAS process. Specifically, the training does not terminate until the loss converges to a near-zero threshold. At this point, the generated circuit is functionally equivalent to the target truth table, ensuring perfect accuracy. This is not merely an empirical observation but a direct result of the rigorous optimization process in DAS, which enforces logical correctness by design. These results validate our hypothesis that integrating learned structural priors with CircuitAR enables more sample-efficient circuit generation compared to basic CircuitNN [14] and template-driven [31] DAS approaches.
Circuit Size. Compared to prior probabilistic generative models [21, 28, 36], our method achieves an order-of-magnitude improvement in directly generatable circuit scale, which is a significant advance, especially given the exponential complexity growth typical in the scaling of circuit size. In practical circuit optimization [13], large circuits are typically partitioned into smaller subcircuits for tractable optimization. Our primary focus is to explore the direct capabilities of generative models in circuit generation. Current evaluation allows us to evaluate the core contributions of our approach without focusing on the additional complexity of decomposition and reintegration.
@@ 4.4 Ablation Study
Figure 8: Ablation study on the initialization with edge probability generated by CircuitAR-2B.
@@
We conducted an ablation study to evaluate the effectiveness of the probability matrix $\hat{\mathcal{A}}$ generated by CircuitAR-2B. As summarized in Figure 8, the experiment results reveal that both variants achieve 100% accuracy across all benchmarks, suggesting that the initialization process does not impair the ability to generate functionally equivalent circuits. The primary distinction lies in the efficiency of the search process, quantified by the number of search steps. Figure 8 underscores the significance of the initialization process, demonstrating that our CircuitAR models can produce high-quality preliminary circuit structures, which can guide the subsequent DAS process effectively.
@@ 5 Conclusion
6 Limitations
This study presents a preliminary investigation of scaling laws under current computational constraints. Due to limited computing resources, the research is intentionally bound to models operating within restricted model parameters and training data. Although our experimental framework demonstrates a tenfold increase in circuit complexity compared to prior works [21, 36, 28], there is substantial potential for further improvement in circuit scale.
Appendix A Related Works
A.1 Autoregressive Modeling
The autoregressive modeling paradigm [24, 27] has been widely adopted for generation tasks in language and vision domains. Built on the transformer architecture, autoregressive models are commonly implemented using causal attention mechanisms in language domains [24], which process data sequentially. However, information does not inherently require sequential processing in vision and graph generation tasks. To address this, researchers have employed bidirectional attention mechanisms for autoregressive modeling [20, 27, 7, 19]. This approach predicts the next token based on previously predicted tokens while allowing unrestricted communication between tokens, thereby relaxing the sequential constraints of traditional autoregressive methods. In this paper, we adopt masked autoregressive modeling for circuit generation, leveraging its ability to provide a global perspective and enhance the modeling of complex dependencies.
A.2 Circuit Generation
In addition to the DAS-based approaches, researchers have also explored next-gate prediction techniques inspired by LLMs for circuit generation. Circuit Transformer [21] predicts the next logic gate using a depth-first traversal and equivalence-preserving decoding. SeaDAG [36] employs a semi-autoregressive diffusion model for DAG generation, maintaining graph structure for precise control. ShortCircuit [28] uses a transformer to generate Boolean expressions from truth tables via next-token prediction. However, these methods are limited by their global view, restricting circuit size and failing to reduce the search space. In contrast, our approach uses global-view masked autoregressive decoding to generate circuits while ensuring logical equivalence and significantly reducing the search space during the DAS process.
Appendix B Implementation Details
B.1 CircuitVQ
The training process of CircuitVQ employs a linear learning rate schedule with the AdamW optimizer set at a learning rate of $2\times 10^{-4}$, a weight decay of 0.1, and a batch size of 2048. The model is fine-tuned for 20 epochs on 8$\times$×A100 GPUs with 80G memory each. Moreover, we use the Graphormer [34] as our CircuitVQ architecture, as mentioned before. Specifically, CircuitVQ comprises 6 encoder layers and 1 decoder layer. The hidden dimension and FFN intermediate dimension are 1152 and 3072, respectively. Additionally, the multi-head attention mechanism employs 32 attention heads. For the vector quantizer component, the codebook dimensionality is set to 4 to improve the codebook utilization, and the codebook size is configured to 8192.
B.2 CircuitAR
The training process of CircuitAR employs a linear learning rate schedule with the AdamW optimizer set at a learning rate of $2\times 10^{-4}$, a weight decay of 0.1, and a batch size of 4096. The model is fine-tuned for 20 epochs on 16$\times$×A100 GPUs with 80G memory each. Moreover, we use the Transformer variant of Llama [9] as our CircuitAR architecture as mentioned before. To form different model sizes, we vary the hidden dimension, FFN intermediate dimension, number of heads and number of layers. We present the details of our CircuitAR architecture configurations in Table 3. For the rest of the hyperparameters, we keep them the same as the standard Llama model.
@@ B.2 CircuitAR
Appendix C Training Datasets
@@
To mitigate data imbalance, a constraint is imposed, limiting each truth table to at most $M$ distinct graph topologies. For truth tables with fewer than $M$ representations, logic synthesis techniques (specifically rewriting algorithms) are applied to generate functionally equivalent subcircuits with distinct topologies. This approach ensures topological diversity while maintaining functional equivalence. Finally, the training datasets with around 400000 circuits (average 200 gates per circuit) are generated using circuits from the ISCAS’85 [6], IWLS’05 [2], and EPFL [3]. $M$ is set to 5 during the generation process. The sizes of PI and PO are capped at 15 each in the training dataset, ensuring manageable truth table sizes while maintaining complexity.
Appendix D Data Augmentation
@@
Appendix E Idle NAND Gates
As shown in Figure 9, all models exhibit a gradual decline in BitsD as the isolated gates proportion increases from 0% to 45%. Large model with 2B parameters demonstrates significantly greater robustness, maintaining BitsD values within a narrow range across varying isolation ratios. In contrast, the small model with 300M parameters shows a more pronounced degradation, with BitsD increasing from 0.5317 to 0.5484 under the same conditions. This disparity highlights the enhanced ability of larger models to efficiently utilize NAND gates for implementing the same truth table. The consistently low BitsD observed in the 2B model underscores its practical utility in predefining search spaces for DAS, offering a notable advantage over smaller models.